# Self-supervised visual learning in the low-data regime: a comparative evaluation

### v1

@@ Abstract
Self-Supervised Learning (SSL) is a valuable and robust training methodology for contemporary Deep Neural Networks (DNNs), enabling unsupervised pretraining on a ‘pretext task’ that does not require ground-truth labels/annotation. This allows efficient representation learning from massive amounts of unlabeled training data, which in turn leads to increased accuracy in a ‘downstream task’ by exploiting supervised transfer learning. Despite the relatively straightforward conceptualization and applicability of SSL, it is not always feasible to collect and/or to utilize very large pretraining datasets, especially when it comes to real-world application settings. In particular, in cases of specialized and domain-specific application scenarios, it may not be achievable or practical to assemble a relevant image pretraining dataset in the order of millions of instances or it could be computationally infeasible to pretrain at this scale, e.g., due to unavailability of sufficient computational resources that SSL methods typically require to produce improved visual analysis results. This situation motivates an investigation on the effectiveness of common SSL pretext tasks, when the pretraining dataset is of relatively limited/constrained size. In this context, this work introduces a taxonomy of modern visual SSL methods, accompanied by detailed explanations and insights regarding the main categories of approaches, and, subsequently, conducts a thorough comparative experimental evaluation in the low-data regime, targeting to identify: a) what is learnt via low-data SSL pretraining, and b) how do different SSL categories behave in such training scenarios. Interestingly, for domain-specific downstream tasks, in-domain low-data SSL pretraining outperforms the common approach of large-scale pretraining on general datasets. Grounded on the obtained results, valuable insights are highlighted regarding the performance of each category of SSL methods, which in turn suggest straightforward future research directions in the field.
@@ 1 Introduction
    A taxonomy of modern SSL methods is introduced, accompanied by detailed explanations and insights regarding each main category of approaches. This classification offers clarity regarding the variety of SSL methods currently available, as well as shed light on the advantages and disadvantages of each algorithmic category.
@@
  • 3.
    Grounded on the obtained results, valuable insights are highlighted regarding the performance of each category of SSL methods (or individual SSL approaches), which in turn suggest straightforward future research directions in the field.
Although SSL pretext tasks can be designed and employed for many different types of data (e.g., timeseries [16], text [17], video [18] [19], audio [20], point clouds [21], or even multimodal data [22] [23]), this article focuses on image analysis for computer vision applications. Moreover, it focuses on generic SSL methods and not ones explicitly designed for specific tasks (e.g., for multi-view clustering [24], product attribute recognition [25], etc.). The remainder of this paper is organized as follows: Section 2 presents a taxonomy of the most common pretext tasks for visual SSL, accompanied by detailed algorithmic explanations for each category. In Section 3, the experimental evaluation methodology and all pretraining or downstream datasets/tasks utilized in this work are detailed. Section 4 illustrates and discusses the obtained results, emphasizing on the extraction of valuable observations and insights. Finally, conclusions about the potential of SSL in the low-data regime are drawn in Section 5.
@@ 2 Pretext Tasks
Properties: Highly sensitive to class imbalance; Target outputs can
be self-assigned through clustering or matched by self-distillation.
@@
Properties: Robust to feature scaling; Commonly used to measure
similarity between representation vectors.
@@
Properties: Demands large dataset size for meaningful results;
Typically used in combination with stochastic image augmentation
functions for generating ‘positive’ images.
@@
Properties: Sensitive to outliers; Used in feature alignment and for
shaping targets in image reconstruction generative pretext tasks.
@@
Properties: Lower sensitivity to outliers;Used for shaping targets in
image reconstruction generative pretext tasks.
2.1 Generative pretext
2.1.1 Rotation Prediction
$\mathcal{L}_{Rot}=\mathcal{L}_{CE}(\mathbf{y},\mathbf{p}),$ (1)
2.1.2 Colorization
$\mathcal{L}_{Col}=\mathcal{L}_{MSE}(\mathbf{P},\mathbf{Y}).$ (2)
2.1.3 Super-resolution
$\mathcal{L}_{SR}=\mathcal{L}_{MSE}(\mathbf{P},\mathbf{Y}).$ (3)
2.1.4 BEiT
The loss function employed for BEiT pretraining is:
6:  Compute $\mathcal{L}_{BEiT}$.
2.1.5 MAE
$\mathcal{L}_{Mas}=\mathcal{L}_{MSE}(\mathbf{P},\mathbf{Y}).$ (5)
4:  Compute loss $\mathcal{L}_{Mas}(\mathbf{P},\mathbf{Y})$.
2.1.6 I-JEPA
[…]
2.2 Contrastive pretext
2.2.1 Early methods
2.2.2 MoCo
2.2.3 PIRL
2.2.4 SimCLR
Algorithm 3 SimCLR pseudocode for a single mini-batch
1:  for each $\mathbf{B}^{i}\in\mathcal{B}$ do
5:  end for
6:  for each $i=1$ to $B$ do
7:     Compute $\mathcal{L}_{NT-Xent}(\mathcal{B})$.
9:  end for7
2.2.5 Barlow Twins
2.3 Self-distillation pretext
2.3.1 BYOL
$\mathcal{L}_{BYOL}=\mathcal{L}_{MSE}(\mathbf{b}_{1},\mathbf{b}_{2}).$ (10)
2.3.2 SimSiam
2.3.3 DINO
Algorithm 4 DINO pretext task pseudocode for a single training image
5:  Compute $\mathcal{L}_{DINO}$.
  • 1.
  • 2.
  • 3.
2.4 Clustering pretext
2.4.1 DeepCluster
Algorithm 5 DeepCluster pretext task pseudocode
1:  Initialize parameters of $f(\cdot)$ (e.g., randomly).
2:  while not converged do
5:     for each mini-batch $\mathcal{B}$ sampled from $\mathcal{X}$ do
6:        Compute $\mathcal{L}_{DeepCluster}$ for each image in $\mathcal{B}$.
8:     end for
9:  end while
  • 1.
  • 2.
  • 3.
  • 4.
2.4.2 SeLa
s.t.:
$\sum_{y=1}^{K}q(y|\mathbf{X}_{i})=1,\forall i\in{1,2,\dots,N},$ (16)
2.4.3 SwAV
[…]
$\mathbf{p}^{v*}_{i}=softmax(\mathbf{C}\mathbf{h}^{v*}_{i}),$ (19)
@@ 2.5 Discussion on the behavior of the SSL pretraining categories
Table 3: A succinct comparison of SSL algorithmic families
Attribute Generative Contrastive Self-distillation Clustering
Prior knowledge (augmentations) Low High High High
Susceptibility to solution collapse No No Possible Possible
Negative sampling No Yes No No
Mini-batch size Flexible Large Moderate Moderate
Representations robustness Not evaluated Moderate High High
@@
With respect to the first point, the majority of SSL methods relies on prior knowledge and exhibits various degrees of robustness to such choices. Essentially, the worst offenders are early generative SSL methods (e.g., rotation prediction, colorization, etc.), which can be nowadays considered outdated. Additionally, contrastive, self-distillation and clustering SSL approaches are heavily dependent on data augmentation transformations, which is a form of prior knowledge about the visual medium and/or the specific desired downstream task. The latter observation stems from the fact that not all downstream tasks are equally invariant to the specific image transformations applied during pretraining; e.g., multi-cropping and rotation may be semantically irrelevant for whole-image classification, but not as much in the case of monocular depth estimation. As a result, pretraining with these common augmentation transformations in order to obtain high-level semantic image representations that are invariant to them may be inappropriate for low-level downstream vision tasks. Among the common SSL algorithms, the only ones that do not rely at all on similar forms of prior knowledge are the image reconstruction generative tasks.
@@
The image reconstruction generative SSL approaches also inherently do not suffer from the danger of trivial solutions via representation collapse, which is theoretically an additional advantage. However, in practice, recent methods in categories that are the most susceptible to this scenario, i.e., self-distillation and SSL via clustering, have converged to clever methodological and/or implementation nuisances for essentially bypassing this issue and avoiding collapse altogether. Mainstream contrastive SSL methods, which inherently do not suffer from collapse issues due to negative sampling, typically come with the cost of requiring very large mini-batch sizes or auxiliary structures (e.g., a large memory bank). A well-known exception is Barlow Twins, since it does not use the contrastive loss and does not employ negative sampling, but in fact DNNs pretrained with it struggle to reach good downstream accuracy without the use of very high-dimensional latent representations and a relatively large mini-batch size.
@@
It has to be noted that it is a self-distillation SSL approach, i.e., DINO, that has exhibited strong zero-shot learning capabilities when combined with a complex ViT, with DINOv2 pretrained on an extremely large-scale dataset reaching the zero-shot performance of competing weakly supervised methods. Essentially, very high model complexity and pretraining dataset size contribute significantly to the performance of DINOv2, but nevertheless self-distillation seems to currently dominate the SSL landscape in computer vision. Even so, research on image reconstruction generative approaches and on SSL via on-line clustering is still on-going.
@@
Regarding robustness analysis, several common SSL-learnt representations have been empirically evaluated against downstream distribution shifts and corruptions in [75]. DINO, a self-distillation method, and SwAV, an approach of SSL via on-line clustering, are found to be the most robust. However, no image reconstruction generative pretext task is considered in that comparative evaluation.
Table 3 succinctly compares the four SSL algorithmic categories presented in this section, excluding the now-outdated early generative methods. The table expresses the prevailing tendencies within the state-of-the-art of each family, with individual methods potentially deviating.
@@ 3 Comparative Evaluation Framework in the Low-Data Regime
The goal of this study is to investigate in depth how the different SSL methodologies behave and what eventual downstream recognition performance is achieved, when pretraining occurs in the low-data regime. This stands in contrast to the currently dominant trend of ever-increasing scale, regarding both the pretraining dataset size and the DNN model complexity in SSL research. As argued in Section 1, low-data pretraining remains highly relevant in cases of domain-specific downstream tasks, where generic ImageNet-1k pretraining-derived or zero-shot representations cannot naturally generalize well. Therefore, a thorough empirical comparative evaluation is designed that incorporates all SSL algorithmic categories outlined in Section 2, in order to identify what can be learnt via SSL in low-data computer vision scenarios. As mentioned in Section 1, ‘low-data’ implies a pretraining dataset with a size in the range of  50k -  300k images, which is significantly smaller than the typical dataset sizes used for SSL pretraining in the literature (e.g., ImageNet-1k).
@@
In the defined evaluation framework, 4 widely used SSL methods from different algorithmic categories are compared under different setups on image classification. The considered approaches are MAE, SimCLR, DINO and DeepClusterV2 that belong to the generative, contrastive, self-distillation and clustering SSL category, respectively. It needs to be highlighted that the 4 aforementioned methods do not necessarily correspond to the 4 best-performing in the literature; however, they are relatively recent, representative of their respective algorithmic category and exhibit performance comparable to the state-of-the-art. The employed neural architecture for each considered SSL method has been selected based on the design choices described in the original paper of each approach; thus, ResNet-50 [9] is used for the cases of SimCLR and DeepClusterV2, while ViT-L/16 [39] is utilized for MAE and DINO. This strategy was preferred because not all SSL methods seem to work equally well with Vision Transformers and each SSL approach has an informally declared ‘default’ architecture. For the baselines pretrained in a supervised manner, both ResNet-50 and ViT-L/16 variants have been employed. As an additional baseline, an off-the-shelf pretrained ViT-L/14 DINOv2 model has also been evaluated. In all cases, linear probing has been employed for downstream finetuning, instead of end-to-end finetuning, due to its widespread usage in the SSL literature and its ability to demonstrate the quality of the representations learnt during pretraining. To this end, each pretrained model is kept frozen and its inferred features are utilized to train a simple downstream linear classifier. In the followings, this linear probing setup is implied wherever finetuning is mentioned.
@@
In order to better structure the wide set of performed experiments and to facilitate the derivation of key/detailed insights, the experiments in the defined experimental framework are divided into the following main classes: a) Main experiments, which are related to whole-image classification tasks; these are further broken down into i) object- and scene-centric (based on the recognition target of the downstream task) and ii) transfer-learning and single-dataset setup (taking into account whether different datasets are used for the pretraining and downstream tasks or not), b) Robustness experiments, which investigate robustness aspects of the various SSL pretraining tasks, and c) Domain-specific experiments, which investigate SSL performance in specialized visual domains that differ significantly from typical/conventional natural RBG images.
@@ 3.1 Main experiments
The main experiments concern whole-image classification tasks and are divided into: a) object- and scene-centric, taking into account the semantic granularity of the recognition target of the downstream task, and b) transfer-learning and single-dataset, considering the dataset setup (i.e., when the pretraining dataset and the downstream dataset for supervised finetuning differ or not, respectively). ImageNet-100666This is a low-data variant of the ImageNet-1k dataset [7], transformed according to [76] and Places-100777This is a low-data variant of the Places-365 dataset [77], transformed to Places-100 according to [78] are selected as the pretraining dataset for the object- and scene-centric case, accordingly. Additionally, the considered downstream datasets for the transfer-learning setup are COCO [79] and STL-10 [80] for the object-centric case, and ADE20K [81] and SUN Database [82] for the scene-centric one. For the single-dataset setup, where both the SSL pretraining and the subsequent supervised downstream finetuning tasks are conducted on a single dataset, ImageNet-100 and Places-100 are used for the object- and the scene-centric cases, respectively. Table 4 summarizes the configurations of the considered main experiments, while notation is also included for clarity purposes.
@@ 3.3 Domain-specific experiments
Specialized image domains differing significantly from conventional/natural RBG images constitute a particularly compelling evaluation scenario for SSL, due to the relative scarcity of (sufficiently) large domain-specific datasets and/or ground-truth annotations. In order to evaluate whether under such evaluation settings domain-specific pretraining in the low-data regime outperforms generic low-data pretraining on ImageNet-100 and/or large-scale pretraining on ImageNet-1k, in terms of downstream accuracy, the selected SSL methods are evaluated under the transfer-learning setup in three different (and significantly diverse) visual domains: a) remote sensing (high altitude aerial photographs), b) medical imaging (i.e., chest X-rays), and c) security imaging (i.e., airport luggage X-rays). The linear probing finetuning protocol for whole-image classification is employed here as well. In all cases, in-domain SSL pretraining is compared against supervised pretraining on ImageNet-100 and on ImageNet-1k, as well as against an off-the-shelf non-low-data-pretrained DINOv2 baseline.
All domain-specific datasets, utilized either for pretraining or downstream finetuning/evaluation, are low-data in the sense defined in this study. For the remote sensing domain, MLRSNet [89] and AID [91] have been used as the pretraining and downstream datasets, respectively, while indicative exemplary images are shown in Figure 3. Regarding the medical imaging domain, the MedPix [90] and ChestX-Det [92] datasets have been used for pretraining and downstream finetuning, respectively, while example images are shown in Figure 4. Moreover, for the security imaging domain, two different subsets of the SIXRay [93] dataset have been used as the pretraining and downstream subset, namely SIXRay-100 and SIXRay-10, respectively, while indicative examples are shown in Figure 5.

### v2

@@ Abstract
Self-Supervised Learning (SSL) is a valuable and robust training methodology for contemporary Deep Neural Networks (DNNs), enabling unsupervised pretraining on a ‘pretext task’ that does not require ground-truth labels/annotation. This allows efficient representation learning from massive amounts of unlabeled training data, which in turn leads to increased accuracy in a ‘downstream task’ by exploiting supervised transfer learning. Despite the relatively straightforward conceptualization and applicability of SSL, it is not always feasible to collect and/or to utilize very large pretraining datasets, especially when it comes to real-world application settings. In particular, in cases of specialized and domain-specific application scenarios, it may not be achievable or practical to assemble a relevant image pretraining dataset in the order of millions of instances or it could be computationally infeasible to pretrain at this scale, e.g., due to unavailability of sufficient computational resources that SSL methods typically require to produce improved visual analysis results. This situation motivates an investigation on the effectiveness of common SSL pretext tasks, when the pretraining dataset is of relatively limited/constrained size. This work briefly introduces the main families of modern visual SSL methods and, subsequently, conducts a thorough comparative experimental evaluation in the low-data regime, targeting to identify: a) what is learnt via low-data SSL pretraining, and b) how do different SSL categories behave in such training scenarios. Interestingly, for domain-specific downstream tasks, in-domain low-data SSL pretraining outperforms the common approach of large-scale pretraining on general datasets. Grounded on the obtained results, valuable insights are highlighted regarding the performance of each category of SSL methods, which in turn suggest straightforward future research directions in the field.
@@ 1 Introduction
    Grounded on the obtained results, valuable insights are highlighted regarding the performance of each main category of SSL methods (or individual SSL approaches), which in turn suggest straightforward future research directions in the field.
Although SSL pretext tasks can be designed and employed for many different types of data (e.g., timeseries [16], text [17], video [18] [19], audio [20], point clouds [21], or even multimodal data [22] [23]), this article focuses on image analysis for computer vision applications. Moreover, it focuses on generic SSL methods and not ones explicitly designed for specific tasks (e.g., for multi-view clustering [24], product attribute recognition [25], etc.). The remainder of this paper is organized as follows: Section 2 briefly presents the most common categories of pretext tasks for visual SSL. In Section 3, the experimental evaluation methodology and all pretraining or downstream datasets/tasks utilized in this work are detailed. Section 4 illustrates and discusses the obtained results, emphasizing on the extraction of valuable observations and insights. Conclusions about the potential of SSL in the low-data regime are drawn in Section 5. Finally, the most important algorithms falling under each of the main SSL pretext task categories are detailed in the Appendix.
@@ 2.1 Main categories of SSL pretext tasks for images
Algorithmic details about a selection of important SSL methods for each of the four categories are provided in the Appendix of this article. Further details and a comprehensive taxonomy can be found in the survey [34].
@@ 2.2 Discussion on the behavior of the SSL pretraining categories
With respect to the first point, the majority of SSL methods relies on prior knowledge and exhibits various degrees of robustness to such choices. Essentially, the worst offenders are early generative SSL methods (e.g., rotation prediction, colorization, jigsaw puzzle reassembly, etc.), which exploit known visual structure of common images [36] and can be nowadays considered outdated. Additionally, contrastive, self-distillation and clustering SSL approaches heavily depend on the data augmentation transformations applied during pretraining, which is a form of prior knowledge about the visual medium [37]. On top of this, proper selection of said image transformations relies on knowledge of the specific desired downstream tasks, due to different invariances. For instance, multi-cropping and rotation may be semantically irrelevant for whole-image classification, but not as much in the case of monocular depth estimation, while aggressive cropping aids occlusion invariance but potentially limits other desirable invariances [38]. As a result, pretraining with these common augmentation transformations in order to obtain high-level semantic image representations that are invariant to them may be inappropriate for low-level downstream vision tasks. Among the common SSL algorithms, the only ones that do not rely at all on similar forms of prior knowledge are the image reconstruction generative tasks [39].
@@
The image reconstruction generative SSL approaches also inherently do not suffer from the danger of trivial solutions via representation collapse, which is theoretically an additional advantage. However, in practice, recent methods in categories that are the most susceptible to this scenario, i.e., self-distillation and SSL via clustering, have converged to clever methodological and/or implementation nuisances for essentially bypassing this issue and avoiding collapse altogether [40]. Mainstream contrastive SSL methods, which inherently do not suffer from collapse issues due to negative sampling, typically come with the cost of requiring very large mini-batch sizes or auxiliary structures (e.g., a large memory bank) [41]. A well-known exception is Barlow Twins [42], since it does not use the contrastive loss and does not employ negative sampling, but in fact DNNs pretrained with it struggle to reach good downstream accuracy without the use of very high-dimensional latent representations [42] [43].
@@
It has to be noted that it is a self-distillation SSL approach, i.e., DINO [44], that has exhibited strong zero-shot learning capabilities when combined with a complex ViT, with DINOv2 [45] pretrained on an extremely large-scale dataset reaching the zero-shot performance of competing weakly supervised methods. Essentially, very high model complexity and pretraining dataset size contribute significantly to the performance of DINOv2, but nevertheless self-distillation seems to currently dominate the SSL landscape in computer vision. Even so, research on image reconstruction generative approaches and on SSL via on-line clustering is still on-going.
@@
Regarding robustness analysis, several common SSL-learnt representations have been empirically evaluated against downstream distribution shifts and corruptions in [46]. DINO, a self-distillation method, and SwAV [47], an approach of SSL via on-line clustering, are found to be the most robust. However, no image reconstruction generative pretext task is considered in that comparative evaluation. On the other hand, the reconstructive method MAE [48] is explicitly tested with regard to its pretraining data efficiency in [49]: the study concludes that MAE can learn occlusion-invariant image representations from rather small datasets, but lacks a systematic comparative evaluation of the issue. Nevertheless, it is shown in [50] that reconstructive generative pretext tasks lead to overfitting on small pretraining datasets, if the underlying DNN model is sufficiently complex.
@@ 3 Comparative Evaluation Framework in the Low-Data Regime
The goal of this study is to investigate in depth how the different SSL methodologies behave and what eventual downstream recognition performance is achieved, when pretraining occurs in the low-data regime. This stands in contrast to the currently dominant trend of ever-increasing scale, regarding both the pretraining dataset size and the DNN model complexity in SSL research. As argued in Section 1, low-data pretraining remains highly relevant in cases of domain-specific downstream tasks, where generic ImageNet-1k pretraining-derived or zero-shot representations cannot naturally generalize well. Therefore, a thorough empirical comparative evaluation is designed that incorporates all SSL algorithmic categories outlined in Section 2, in order to identify what can be learnt via SSL in low-data computer vision scenarios.
As mentioned in Section 1, ‘low-data’ in this article implies a pretraining dataset with a total size in the range of  50k -  300k images, which is significantly smaller than the typical dataset sizes used for SSL pretraining in the literature (e.g., ImageNet-1k). Early exploratory experiments indicated that SSL pretraining is largely ineffective on datasets with a total size below this range, i.e., lower than  50k images, since the lack of ground-truth labels makes it critical to have enough images for learning adequate representations. This observation is in line with published literature (e.g., [50] [51]), leading to a reasonable choice of  50k images (in total) as the lower bound of a meaningful SSL low-data regime. Still, even in such datasets, their actual training subset is obviously in fact smaller.
@@
In the defined evaluation framework, 4 widely used SSL methods from different algorithmic categories are compared under different setups on image classification and, in one case, on object detection. The considered approaches are MAE [48], SimCLR [52], DINO [44] and DeepClusterV2 [47] that belong to the generative, contrastive, self-distillation and clustering SSL category, respectively. It needs to be highlighted that the 4 aforementioned methods do not necessarily correspond to the 4 best-performing in the literature; however, they are relatively recent, representative of their respective algorithmic category and exhibit performance comparable to the state-of-the-art. The employed neural architecture for each considered SSL method has been selected based on the design choices described in the original paper of each approach; thus, ResNet-50 [9] is used for the cases of SimCLR and DeepClusterV2, while ViT-L/16 [53] is utilized for MAE and DINO. This strategy was preferred because not all SSL methods seem to work equally well with Vision Transformers and each SSL approach has an informally declared ‘default’ architecture. For the baselines pretrained in a supervised manner, both ResNet-50 and ViT-L/16 variants have been employed. As an additional baseline, an off-the-shelf pretrained ViT-L/14 DINOv2 model has also been evaluated. In all cases, except in the object detection experiment, linear probing has been employed for downstream finetuning, instead of end-to-end finetuning, due to its widespread usage in the SSL literature and its ability to demonstrate the quality of the representations learnt during pretraining. To this end, each pretrained model is kept frozen and its inferred features are utilized to train a simple downstream linear classifier. In the followings, this linear probing setup is implied wherever finetuning is mentioned for whole-image classification.
@@
In order to better structure the wide set of performed experiments and to facilitate the derivation of key/detailed insights, the experiments in the defined experimental framework are divided into the following main classes: a) Main experiments, which are related to whole-image classification tasks; these are further broken down into i) object- and scene-centric (based on the recognition target of the downstream task) and ii) transfer-learning and single-dataset setup (taking into account whether different datasets are used for the pretraining and downstream tasks or not), b) Robustness experiments, which investigate robustness aspects of the various SSL pretraining tasks, and c) Domain-specific experiments, which investigate SSL performance in specialized visual domains that differ significantly from typical/conventional natural RBG images. Four different domains are evaluated using a whole-image classification downstream task, while one domain (UAV-captured infrared footage) is evaluated on an object detection downstream task.
@@ 3.1 Main experiments
The main experiments concern whole-image classification tasks and are divided into: a) object- and scene-centric, taking into account the semantic granularity of the recognition target of the downstream task, and b) transfer-learning and single-dataset, considering the dataset setup (i.e., when the pretraining dataset and the downstream dataset for supervised finetuning differ or not, respectively). ImageNet-100222This is a low-data variant of the ImageNet-1k dataset [7], transformed according to [54] and Places-100333This is a low-data variant of the Places-365 dataset [55], transformed to Places-100 according to [56] are selected as the pretraining dataset for the object- and scene-centric case, accordingly. Additionally, the considered downstream datasets for the transfer-learning setup are COCO [57] and STL-10 [58] for the object-centric case, and ADE20K [59] and SUN Database [60] for the scene-centric one. For the single-dataset setup, where both the SSL pretraining and the subsequent supervised downstream finetuning tasks are conducted on a single dataset, ImageNet-100 and Places-100 are used for the object- and the scene-centric cases, respectively. Table 3 summarizes the configurations of the considered main experiments, while notation is also included for clarity purposes.
@@ 3.3 Domain-specific experiments
Figure 6: Example images from the unlabeled CEM500K microscopy imaging dataset [70], depicting a) brain cells of drosophila tissue, and b) brain cells of human tissue
Figure 7: Example images from the object detection UAV-captured infrared dataset [70], depicting objects of classes a) ‘Animal’, and b) ‘Human’
Specialized image domains differing significantly from conventional/natural RBG images constitute a particularly compelling evaluation scenario for SSL, due to the relative scarcity of (sufficiently) large domain-specific datasets and/or ground-truth annotations. In order to evaluate whether under such evaluation settings domain-specific pretraining in the low-data regime outperforms generic low-data pretraining on ImageNet-100 and/or large-scale pretraining on ImageNet-1k, in terms of downstream accuracy, the selected SSL methods are evaluated under the transfer-learning setup in five different (and significantly diverse) visual domains: a) remote sensing (high altitude aerial photographs), b) medical imaging (i.e., chest X-rays), c) security imaging (i.e., airport luggage X-rays), d) microscopy imaging (e.g., blood cells), and e) UAV-captured infrared footage (i.e., aerial thermal images). Evaluation in the first four domains is conducted on a whole-image classification downstream task, while in the UAV-captured infrared domain on an object detection downstream task.
For whole-image classification, the linear probing finetuning protocol classification is employed here as well. For the object detection experiment, the pretrained backbones are combined with the recent real-time object detector YOLOv8 [71]. During finetuning, each backbone is kept frozen so that only the neck and detection head modules of YOLOv8 are optimized. In all cases, in-domain SSL pretraining is compared against supervised pretraining on ImageNet-100 and on ImageNet-1k, as well as against an off-the-shelf non-low-data-pretrained DINOv2 baseline.
For the remote sensing domain, MLRSNet [67] and AID [72] have been used as the pretraining and downstream datasets, respectively, while indicative exemplary images are shown in Figure 3. Regarding the medical imaging domain, the MedPix [68] and ChestX-Det [73] datasets have been used for pretraining and downstream finetuning, respectively, while example images are shown in Figure 4. Moreover, for the security imaging domain, two different subsets of the SIXRay [69] dataset have been used as the pretraining and downstream subset, namely SIXRay-100 and SIXRay-10, respectively, while indicative examples are shown in Figure 5. For the microscopy imaging domain, CEM500K [70] and RaabinWBC [74] have been employed as pretraining and downstream datasets, respectively, while example images are depicted in Figure 6. Finally, for the UAV-captured infrared domain, BIRDSAI [75] and HIT-UAV [76] have been employed as pretraining and downstream datasets, respectively, with indicative images shown in Figure 7. All domain-specific datasets, utilized either for pretraining or downstream finetuning/evaluation, are low-data in the sense defined in this study (after appropriate subsampling, in the case of CEM500K).